# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 4.1. Sheaves on $\mathsf{ED}$ spaces and their products
For the base case $n=1$, the functions $\tau_{\alpha}$ appearing in the proof of item 5 of Theorem 2.10 furnish a nontrivial coherent $\Phi=\langle\varphi_{\alpha}:\kappa\times\alpha\to K\mid\alpha\in\omega_{1}\rangle$ with $\kappa=\aleph_{0}$, as the reader may verify. In order to better fix both our notation and a template for higher $n$, though, we will record a more hands-on construction of such a $\Phi$, one in which $\kappa=\omega_{1}$. For this purpose, fix a sequence $\langle C_{\alpha}\subseteq\alpha\mid\alpha\in\mathrm{Cof}(\omega)\cap\omega_{1}\rangle$ in which $\mathrm{otp}(C_{\alpha})=\omega$ and $\mathrm{sup}(C_{\alpha})=\alpha$ for every countable limit ordinal $\alpha$; for each such $\alpha$, let $\chi_{\alpha}(0,-):\alpha\to K$ be the characteristic function of $C_{\alpha}$ and let $\chi_{\alpha}:\omega_{1}\times\alpha\to K$ otherwise equal 00. For any $\psi:\kappa\times\alpha\to K$ and $\gamma<\kappa$, let $\psi^{(\gamma)}(\gamma+\eta,\xi)=\psi(\eta,\xi)$ for all $(\eta,\xi)\in\kappa\times\alpha$ and let $\psi^{(\gamma)}$ otherwise equal 00. We proceed now to construct by recursion on $\alpha$ a nontrivial coherent $\Phi=\langle\varphi_{\alpha}:\omega_{1}\times\alpha\to K\mid\alpha\in\omega_{1}\rangle$, as desired.
@@
    $\beta=\alpha+1$ for some $\alpha$. In this case, let $\varphi_{\beta}(\eta,\xi)=\varphi_{\alpha}(\eta,\xi)$ for $\xi<\alpha$ and otherwise equal 00.
@@
for some $\xi<\alpha_{0}$ (if not, then $\Upsilon\restriction(\beta)^{n}$ would itself be trivial on the domain $[\beta,\beta+\omega_{n})\times\beta$, implying that the associated $\Psi$ is). By Fodor’s Lemma, $f$ is constantly $\vec{\alpha}$ on some cofinal $B\subseteq\omega_{n+1}$, hence the witnesses $\eta$ to equation 17 for this $\vec{\alpha}$ are cofinal in $\omega_{n+1}$ as well. Since $\mathrm{supp}(\varphi_{\vec{\alpha}})\subseteq\mathrm{min}\,B$, though, this implies that $\Upsilon$ does not trivialize $\Phi$. This contradiction completes the argument. ∎
@@
  • by the discussion preceding Theorem 4.10, $\kappa$ must be infinite;
@@
  • it is consistent with the $\mathsf{ZFC}$ axioms that $\kappa=\aleph_{0}$ suffices for all $n>0$ (it suffices in the constructible universe $L$, by straightforward modifications of [15, §3.3], for example);
  • but whether $\kappa=\aleph_{0}$ suffices for all $n>0$ in $\mathsf{ZFC}$ alone is a subtle question, one again closely related to that of the groups $\mathrm{H}^{n}(\omega_{n};\mathbb{Z})$ referenced in Section 3.
@@
If $S=\beta X$ for a discrete space $X$ with $|P(X)|<\aleph_{\omega}$ then for any field $K$ and $K$-vector space $W$, the constant sheaf $\mathcal{W}$ on $S$ is of finite injective dimension.
@@
Let $|P(X)|=\aleph_{n}$; any $\mathcal{I}_{U}$ as in equation 16 above then determines a $\mathrm{Cof}(\leq\aleph_{n})$-indexed inverse system with surjective bonding maps whose $\mathrm{Rlim}$ computes $\mathrm{H}^{q}(U;\mathcal{W})$. By equations 13 through 16 together with Goblot’s Theorem, the second, and hence the first, of Lemma 4.8’s conditions therefore holds. ∎
@@
If $\aleph_{\omega}$ is a strong limit cardinal then for any field $K$ and $K$-vector space $W$ and $\mathsf{ED}$ space $S$ of cardinality less than $\aleph_{\omega}$, the constant sheaf $\mathcal{W}$ on $S$ is of finite injective dimension.
@@
Just before Theorem 4.10, we noted that for any field $K$ and $\mathsf{ED}$ space $S$ the injective dimension of the sheaf $\mathcal{K}$ on $S$ is at most one. The following refinement completes the proof of Theorem 4.3.
Lemma 4.14.
For any finite field $K$ and $\mathsf{ED}$ space $S$, the sheaf $\mathcal{K}$ on $S$ is injective.
Proof.
Observe first that since $K$ is compact, any continuous map from an open $U\subseteq S$ to $K$ uniquely extends to a continuous map $\beta U\to K$. Similarly, since $S$ is $\mathsf{ED}$, the inclusion $U\hookrightarrow S$ uniquely extends to an embedding of $\beta U$ as an open and closed subspace of $S$. Together these facts imply that the restriction map $\mathcal{K}(S)\to\mathcal{K}(U)$ is surjective. When the latter holds, as here, for all open $U\subseteq S$, we say that the sheaf in question is flasque (the term’s application to inverse systems above was a special instance); the interest of this property is that flasque resolutions are acyclic and therefore convenient for cohomology computations. It is easy to see also that $\iota^{*}$ preserves this property for any such $\iota:U\hookrightarrow S$. Putting these facts together, we have that $0=\mathrm{H}^{1}(U;\iota^{*}\mathcal{K})=\mathrm{Ext}^{1}(\iota_{!}\mathcal{K},\mathcal{K})$ for all such $\iota$, and conclude, by Claim 4.9, that $\mathcal{K}$ is injective. ∎
@@ 4.2. Products of compact projective anima
Taken together, an object $X\in\mathcal{C}$ is compact projective if $\mathrm{Hom}(X,-)$ commutes with all filtered colimits and reflexive coequalizers: equivalently, it commutes with all (so-called) $1$-sifted colimits. [73, p. 74]
This suffices, for our purposes, to describe the latter class of colimits, although a more intuitive characterization is the following: filtered colimits are precisely those which commute with finite limits; (1(1( 1-)sifted colimits are precisely those which commute with finite products. The prefix $1$-, of course, signals that we’re discussing the ordinary category theoretic version of what we’ll soon upgrade to an $\infty$-category theoretic notion. Write $\mathsf{sInd}(\mathcal{C})$ for the free cocompletion of $\mathcal{C}$ with respect to $1$-sifted colimits. If $\mathcal{C}$ is cocomplete then we have a fully faithful embedding $\mathsf{sInd}(\mathcal{C}^{\mathrm{cp}})\to\mathcal{C}:\textnormal{``}\mathrm{colim}\textnormal{''}\,X_{i}\mapsto\mathrm{colim}\,X_{i}$, and “if $\mathcal{C}$ is generated under small colimits by $\mathcal{C}^{\mathrm{cp}}$ then that functor is an equivalence” ([73, p. 74]; for fuller argument, see [81, §5.1.1]). In particular, we have $\mathsf{sInd}(\mathcal{C}^{\mathrm{cp}})\cong\mathcal{C}$ for each of the enumerated examples of the section’s introduction: the category $\mathsf{Set}$ is the $1$-sifted ind-completion $\mathsf{sInd}(\mathsf{Set}^{\mathrm{cp}})$ of the category $\mathsf{Set}^{\mathrm{cp}}$ of finite sets; similarly for $\mathsf{Ab}$, for $\mathsf{Cond(Ab)}$, and so on.
Animation is simply the $\infty$-category theoretic version of this operation; [^13] For sifted diagrams in $\infty$-categories, see [50, §5.5.8] or [52, Tag 02QD]; these are again, intuitively, those whose colimits preserve finite products. continuing with [73], we have:
@@
    The animation $\mathsf{Ani(Set)}$ of the category of sets is nothing other than the $\infty$-category of “spaces”: it is, in other words, that higher category theoretic incarnation of the homotopy category of topological spaces playing within $\infty$-category theory a role broadly analogous to that of $\mathsf{Set}$ within ordinary category theory (see [50, §1.2.16] for other presentations). Its objects are termed anima, and the category itself is frequently abbreviated $\mathsf{Ani}$.
@@
Similarly for the categories $\mathsf{Cond(Set)}$ and $\mathsf{Cond(Ab)}$; for these and the above examples and the following definition and lemma, see again [73, §11], where the terminology nonabelian derived category for the animation construction is also noted. As it happens, the $\infty$-categorical condensation and animation operations commute:
@@
[…]
in $\mathcal{C}$ with $X$ in the degree $-1$. (Above, we’ve left degeneracy maps from lower to higher degrees undepicted simply for typographical reasons, and we will tend in what follows to denote such a hypercover simply by $R_{\bullet}$, i.e., to identify it with its non-negatively graded portion, so that a more scrupulous notation would be $R_{\bullet}\to X$.) Like covers or resolutions, hypercovers facilitate cohomology computations by decomposing their target into more convenient objects, and the induced hypercovering system of higher-order intersections $\bigcap_{i\in I}U_{i}$ of elements of an open cover $\mathcal{U}$ of a topological space $X$ which underlies the Čech cochain complex of $(X,\mathcal{U})$ is among the most fundamental of examples. The more general notion of a hypercover is often motivated as allowing for further refinements of the cover at each level $n$ of the complex, a prospect technically managed by an $n$-coskeleton condition for which we refer the reader to any standard reference ([1, §8], [31, §4], [26, §4]). The important points for our purposes are the following:
@@
    In the present context, the “convenient objects” are the $\mathsf{ED}$ spaces; in particular, any $X\in\mathsf{CHaus}$ admits a hypercovering $R_{\bullet}$ by $\mathsf{ED}$ spaces, and even a standard such “free resolution” [10, 2.2.5]. For hypercoverings of topological spaces $R_{\bullet}\to X$, we have moreover that $X$ is weakly homotopy equivalent to the homotopy colimit of $R_{\bullet}$ [32].
@@
    Just as was the case for ordinary categories (Definition 2.1), the sheaf condition defining condensed $\infty$-categories considerably simplifies when we restrict the underlying site to $\mathsf{ED}$ spaces, a simplicity on display in Definition 4.16. In the $\infty$-categorical setting, though, that definition is equivalent not to the category of sheaves on the profinite site, but of hypersheaves (see [24, §9], [10, §2.2.2–2.2.7] for this point in the pyknotic setting, and [43, §A.4] for a concise review of the relevant notions). In other words, writing $\mathsf{StrLim}$ for the class of strong limit cardinals, a condensed anima may equivalently be characterized as an element of $\mathrm{colim}_{\kappa\in\mathsf{StrLim}}\,\mathrm{Fun}(\mathsf{ProFin}_{\kappa}^{\mathrm{op}},\mathsf{Ani})$ satisfying the $\infty$-categorical analogues of conditions (1) through (3) of Definition 2.1; the equalizer condition
@@
    for any hypercover $T_{\bullet}$ of $S$ in the category of profinite spaces (cf. [43, A.4.20] and the references therein). Note that this condition in turn implies that $S=\mathrm{colim}\,T_{\bullet}$ when each is interpreted in the category $\mathsf{Cond(Ani)}$. This follows from the Yoneda lemma, together with the fact that
@@
For all fields $K$ and $\mathsf{ED}$ spaces $S$ and $T$, if $S\times T$, viewed as a product of condensed anima, is compact, then the sheaf $\mathcal{K}$ on the topological space $S\times T$ is of finite injective dimension.
@@
For the second equality, recall that $\mathrm{RHom}$ preserves $\infty$-categorical limits in the second coordinate; the third is just the pushforward-pullback adjunction. When $\mathcal{B}=\mathcal{K}$, though, we know from Theorem 4.3 that $f_{i}^{*}\mathcal{B}$ is of injective dimension at most $1$. Hence as $\mathcal{A}$ ranges over sheaves of $K$-modules on $S\times T$, the complex $\mathrm{RHom}(\mathcal{A},g_{*}g^{*}\mathcal{K})$ is a finite limit of uniformly bounded complexes and therefore itself uniformly bounded. So too, in consequence, are its retracts $\mathrm{RHom}(\mathcal{A},\mathcal{K})$, a conclusion amounting precisely to our assertion that the sheaf $\mathcal{K}$ on $S\times T$ is of finite injective dimension.
It remains only to justify our claim; in essence, it’s the sum of two subclaims. The first is that the functor $D$ which sends each profinite set $T$ to the $\infty$-category of sheaves on $T$, and morphism $f:S\to T$ to $f^{*}:D(T)\to D(S)$, satisfies descent, i.e., satisfies the condition (18) for the functor $F=D$ [43, Lem. 3.5.12]. The second is that coupling this first subclaim with Lemma D.4.7(ii) of [43] implies the claim: within the lemma, let $I$ denote the diagram $\Delta_{\leq n}^{\mathrm{op}}$ defining $X_{n}$, let $\mathcal{D}_{i}=D(R_{i})$, let $\mathcal{C}=D(S\times T)$, let $F_{i}=f_{i}^{*}$, and let $F=g^{*}$. Then by descent, $D(X_{n})=\mathrm{lim}_{I}\,D(R_{i})$, and the lemma defines the right adjoint $g^{*}$ of $g_{*}$ as a limit of the right adjoints $f_{i^{*}}$ of $f_{i}^{*}$ in the manner, precisely, of our claim. ∎

### v2

@@ 4.1. Sheaves on $\mathsf{ED}$ spaces and their products
For the base case $n=1$, the functions $\tau_{\alpha}$ appearing in the proof of item 5 of Theorem 2.10 furnish a nontrivial coherent $\Phi=\langle\varphi_{\alpha}:\kappa\times\alpha\to K\mid\alpha\in\omega_{1}\rangle$ with $\kappa=\aleph_{0}$, as the reader may verify. In order to better fix both our notation and a template for higher $n$, though, we will record a more hands-on construction of such a $\Phi$, one in which $\kappa=\omega_{1}$. For this purpose, fix a sequence $\langle C_{\alpha}\subseteq\alpha\mid\alpha\in\mathrm{Cof}(\omega)\cap\omega_{1}\rangle$ in which $\mathrm{otp}(C_{\alpha})=\omega$ and $\mathrm{sup}(C_{\alpha})=\alpha$ for every countable limit ordinal $\alpha$; for each such $\alpha$, let $\chi_{\alpha}(0,-):\alpha\to K$ be the characteristic function of $C_{\alpha}$ and let $\chi_{\alpha}:\omega_{1}\times\alpha\to K$ otherwise equal 0. For any $\psi:\kappa\times\alpha\to K$ and $\gamma<\kappa$, define $\psi^{(\gamma)}:\kappa\times\alpha\rightarrow K$ by letting $\psi^{(\gamma)}(\gamma+\eta,\xi)=\psi(\eta,\xi)$ for all $(\eta,\xi)\in\kappa\times\alpha$ and letting $\psi^{(\gamma)}$ otherwise equal 0. We proceed now to construct by recursion on $\alpha$ a nontrivial coherent $\Phi=\langle\varphi_{\alpha}:\omega_{1}\times\alpha\to K\mid\alpha\in\omega_{1}\rangle$, as desired.
@@
    $\beta=\alpha+1$ for some $\alpha$. In this case, let $\varphi_{\beta}(\eta,\xi)=\varphi_{\alpha}(\eta,\xi)$ for $\xi<\alpha$ and otherwise equal 0.
@@
for some $\xi<\alpha_{0}$ (if not, then $\Upsilon\restriction(\beta)^{n}$ would itself be trivial on the domain $\beta,\beta+\omega_{n})\times\beta$, implying that the associated $\Psi$ is). By Fodor’s Lemma, $f$ is constantly $\vec{\alpha}$ on some cofinal $B\subseteq\omega_{n+1}$, hence the witnesses $\eta$ to equation 20 for this $\vec{\alpha}$ are cofinal in $\omega_{n+1}$ as well. Since $\mathrm{supp}(\varphi_{\vec{\alpha}})\subseteq\mathrm{min}\,B$, though, this implies that $\Upsilon$ does not trivialize $\Phi$. This contradiction completes the argument. ∎
@@
  • $\kappa$ must be infinite (if it is finite then the system of equation 19 is flasque);
@@
  • it is consistent with the $\mathsf{ZFC}$ axioms that $\kappa=\aleph_{0}$ suffices for all $n>0$ (it suffices in the constructible universe $L$, by straightforward modifications of [17, §3.3], for example);
  • but whether $\kappa=\aleph_{0}$ suffices for all $n>0$ in $\mathsf{ZFC}$ alone is a subtle question, one again closely related to that of the groups $\mathrm{H}^{n}(\omega_{n};\mathbb{Z})$ (or $\mathrm{H}^{n}(\omega_{n};\mathbb{Z}/p\mathbb{Z})$) referenced in Section 3.
@@
If $S=\beta X$ for a discrete space $X$ with $|P(X)|<\aleph_{\omega}$ then for any finite field $K$ and $K$-vector space $W$, the constant sheaf $\mathcal{W}$ on $S$ is of finite injective dimension.
@@
Let $|P(X)|=\aleph_{n}$; any $\mathcal{I}_{U}$ as in equation 19 above then determines a $\mathrm{Cof}(\leq\aleph_{n})$-indexed inverse system with surjective bonding maps whose $\mathrm{Rlim}$ computes $\mathrm{H}^{q}(U;\mathcal{W})$. By equations 16 through 19 together with Goblot’s Theorem, the second, and hence the first, of Lemma 4.10’s conditions therefore holds. ∎
@@
If $\aleph_{\omega}$ is a strong limit cardinal then for any finite field $K$ and $K$-vector space $W$ and $\mathsf{ED}$ space $S$ of cardinality less than $\aleph_{\omega}$, the constant sheaf $\mathcal{W}$ on $S$ is of finite injective dimension.
@@ 4.2. Products of compact projective anima
Taken together, an object $X\in\mathcal{C}$ is compact projective if $\mathrm{Hom}(X,-)$ commutes with all filtered colimits and reflexive coequalizers: equivalently, it commutes with all (so-called) $1$-sifted colimits. [76, p. 74]
Here, filtered colimits are precisely those which commute with finite limits, and $(1\text{-})$sifted colimits are precisely those which commute with finite products. The prefix $1$-, of course, signals that we’re discussing the ordinary category theoretic version of what we’ll soon upgrade to an $\infty$-category theoretic notion. Write $\mathsf{sInd}(\mathcal{C})$ for the free cocompletion of $\mathcal{C}$ with respect to $1$-sifted colimits. If $\mathcal{C}$ is cocomplete then we have a fully faithful embedding $\mathsf{sInd}(\mathcal{C}^{\mathrm{cp}})\to\mathcal{C}:\textnormal{``}\mathrm{colim}\textnormal{''}\,X_{i}\mapsto\mathrm{colim}\,X_{i}$, and “if $\mathcal{C}$ is generated under small colimits by $\mathcal{C}^{\mathrm{cp}}$ then that functor is an equivalence” ([76, p. 74]; for fuller argument, see [84, §5.1.1]). In particular, we have $\mathsf{sInd}(\mathcal{C}^{\mathrm{cp}})\cong\mathcal{C}$ for each of the enumerated examples of the section’s introduction: the category $\mathsf{Set}$ is the $1$-sifted ind-completion $\mathsf{sInd}(\mathsf{Set}^{\mathrm{cp}})$ of the category $\mathsf{Set}^{\mathrm{cp}}$ of finite sets; similarly for $\mathsf{Ab}$, for $\mathsf{Cond(Ab)}$, and so on.
Animation is simply the $\infty$-category theoretic version of this operation; [^13] For sifted diagrams in $\infty$-categories, see [53, §5.5.8] or [55, Tag 02QD]; these are again, intuitively, those whose colimits preserve finite products. continuing with [76], we have:
@@
    The animation $\mathsf{Ani(Set)}$ of the category of sets is nothing other than the $\infty$-category of “spaces”: it is, in other words, that higher category theoretic incarnation of the homotopy category of topological spaces playing within $\infty$-category theory a role broadly analogous to that of $\mathsf{Set}$ within ordinary category theory (see [53, §1.2.16] for other presentations). Its objects are termed anima, and the category itself is frequently abbreviated $\mathsf{Ani}$.
@@
Similarly for the categories $\mathsf{Cond(Set)}$ and $\mathsf{Cond(Ab)}$; for these and the above examples and the following definition and lemma, see again [76, §11], where the terminology nonabelian derived category for the animation construction is also noted. As it happens, the $\infty$-categorical condensation and animation operations commute:
@@
$\cdots\begin{subarray}{c}\longrightarrow\\[-10.00002pt] \longrightarrow\\[-10.00002pt] \longrightarrow\end{subarray}R_{1}\begin{subarray}{c}\longrightarrow\\[-10.00002pt] \longrightarrow\end{subarray}R_{0}\begin{subarray}{c}\longrightarrow\end{subarray}X$
in $\mathcal{C}$ with $X$ in the degree 1-1. (Above, we’ve left degeneracy maps from lower to higher degrees undepicted simply for typographical reasons, and we will tend in what follows to denote such a hypercover simply by $R_{\bullet}$, i.e., to identify it with its non-negatively graded portion, so that a more scrupulous notation would be $R_{\bullet}\to X$.) Like covers or resolutions, hypercovers facilitate cohomology computations by decomposing their target into more convenient objects, and the induced hypercovering system of higher-order intersections $\bigcap_{i\in I}U_{i}$ of elements of an open cover $\mathcal{U}$ of a topological space $X$ which underlies the Čech cochain complex of $(X,\mathcal{U})$ is among the most fundamental of examples. The more general notion of a hypercover is often motivated as allowing for further refinements of the cover at each level $n$ of the complex, a prospect technically managed by an $n$-coskeleton condition for which we refer the reader to any standard reference ([1, §8], [33, §4], [28, §4]). The important points for our purposes are the following:
@@
    In the present context, the “convenient objects” are the $\mathsf{ED}$ spaces; in particular, any $X\in\mathsf{CHaus}$ admits a hypercovering $R_{\bullet}$ by $\mathsf{ED}$ spaces, and even a standard such “free resolution” [11, 2.2.5]. For hypercoverings of topological spaces $R_{\bullet}\to X$, we have moreover that $X$ is weakly homotopy equivalent to the homotopy colimit of $R_{\bullet}$ [34].
@@
    Just as was the case for ordinary categories (Definition 2.1), the sheaf condition defining condensed $\infty$-categories considerably simplifies when we restrict the underlying site to $\mathsf{ED}$ spaces, a simplicity on display in Definition 4.16. In the $\infty$-categorical setting, though, that definition is equivalent not to the category of sheaves on the profinite site, but of hypersheaves (see [26, §9], [11, §2.2.2–2.2.7] for this point in the pyknotic setting, and [46, §A.4] for a concise review of the relevant notions). In other words, writing $\mathsf{StrLim}$ for the class of strong limit cardinals, a condensed anima may equivalently be characterized as an element of $\mathrm{colim}_{\kappa\in\mathsf{StrLim}}\,\mathrm{Fun}(\mathsf{ProFin}_{\kappa}^{\mathrm{op}},\mathsf{Ani})$ satisfying the $\infty$-categorical analogues of conditions (1) through (3) of Definition 2.1; the equalizer condition
@@
    for any hypercover $T_{\bullet}$ of $S$ in the category of profinite spaces (cf. [46, A.4.20] and the references therein). Note that this condition in turn implies that $S=\mathrm{colim}\,T_{\bullet}$ when each is interpreted in the category $\mathsf{Cond(Ani)}$. This follows from the Yoneda lemma, together with the fact that
@@
For all finite fields $K$ and $\mathsf{ED}$ spaces $S$ and $T$, if $S\times T$, viewed as a product of condensed anima, is compact, then the sheaf $\mathcal{K}$ on the topological space $S\times T$ is of finite injective dimension.
@@
For the second equality, recall that $\mathrm{RHom}$ preserves $\infty$-categorical limits in the second coordinate; the third is just the pushforward-pullback adjunction. When $\mathcal{B}=\mathcal{K}$, though, we know from Theorem 4.3 that $f_{i}^{*}\mathcal{B}$ is injective. Hence as $\mathcal{A}$ ranges over sheaves of $K$-modules on $S\times T$, the complex $\mathrm{RHom}(\mathcal{A},g_{*}g^{*}\mathcal{K})$ is a finite limit of uniformly bounded complexes and therefore itself uniformly bounded. So too, in consequence, are its retracts $\mathrm{RHom}(\mathcal{A},\mathcal{K})$, a conclusion amounting precisely to our assertion that the sheaf $\mathcal{K}$ on $S\times T$ is of finite injective dimension.
It remains only to justify our claim; in essence, it’s the sum of two subclaims. The first is that the functor $D$ which sends each profinite set $T$ to the $\infty$-category of sheaves on $T$, and morphism $f:S\to T$ to $f^{*}:D(T)\to D(S)$, satisfies descent, i.e., satisfies the condition (21) for the functor $F=D$ [46, Lem. 3.5.12]. The second is that coupling this first subclaim with Lemma D.4.7(ii) of [46] implies the claim: within the lemma, let $I$ denote the diagram $\Delta_{\leq n}^{\mathrm{op}}$ defining $X_{n}$, let $\mathcal{D}_{i}=D(R_{i})$, let $\mathcal{C}=D(S\times T)$, let $F_{i}=f_{i}^{*}$, and let $F=g^{*}$. Then by descent, $D(X_{n})=\mathrm{lim}_{I}\,D(R_{i})$, and the lemma defines the right adjoint $g^{*}$ of $g_{*}$ as a limit of the right adjoints $f_{i^{*}}$ of $f_{i}^{*}$ in the manner, precisely, of our claim. ∎